OpenStudy (anonymous):

If h(x) = −5x − 10, find h(−2). −20 1.6 −1 0

OpenStudy (anonymous):

h(-2) = -5(-2) - 10 = ?

OpenStudy (anonymous):

10-10 = 0

OpenStudy (anonymous):

\[h \left( x \right)=-5x-10\] put x=-2 \[h \left( -2 \right)=-5\times-2-10=10-10=0\]

OpenStudy (anonymous):

If h(x) = −5x − 10, find h(−2). --SUBSTITUTE h(-2) for x variables. h(-2) = −5(-2) − 10 h(-2) = 10 - 10 h(-2) = 0